Problem 12.118 Part A What osmotic pressure (in atmospheres) would you expect for each of the following solutions? 4.71 g of NaCl in 345.0 mL of aqueous solution at 42°C. atm osm Submit X Incorrect; Try Again; 4 attempts remaining Part B 5.63 g of sodium acetate, CH3 CO2Na, in 52.0 mL of aqueous solution at 20 C atm Submit Previous Answers Request Answer
Answer #1

Part A: Osmotic pressure = MRT(M is mol/lit, R =0.0821 lit.atm K-1&T is absolute temperature)

Osmotic pressure = (4.71/58.4)0.345x 0.0821x315= 0.719 atm.

Part B :

Osmotic pressure =(5.63/82)0.052x0.0821x293=0.085 atm